OpenStudy (anonymous):

How do you solve for x in this equation? 49^x = 7^(x – 3)

OpenStudy (anonymous):

take the natural log

OpenStudy (anonymous):

7^2x=7^(x-3) 2x=x-3

OpenStudy (praxer):

so, x=-3

OpenStudy (anonymous):

Thank you guys so much :)